can you please help me solve this equation:-12=5x-7
negitive twelve equals five x minus seven

Answer by wuwei96815(245)

-12=5x-7
Add 7 to each side of the equation.
-12+7=5x-7+7
-5=5x
Divide each side by 5.
-5/5=5x/5
-1=x